OpenStudy (anonymous):

Find the indicated term of the given arithmetic sequence. a^14 for 320, 316, 312, ... a) 359 b)268 c)264 d)372

OpenStudy (tkhunny):

You have a1. It is 320 You have d. It is 316 - 320 = -4 and 312 - 316 = -4. Go!

OpenStudy (anonymous):

n = 58 \[a_{1} = 320\] d = -4

OpenStudy (anonymous):

\[a _{14}= 320 + (14 - 1)-4\] \[a _{14}= 268\]